OpenStudy (camper4834):

Somebody else get in on this argument... On the graph is point P (1,-2) The center of dilation is at (10,10) The scale factor is 2.5 I say that the answer for the new P also called P' is at (-12.5,-20) The other person says (12.5,5) Someone please say who is correct

OpenStudy (camper4834):

@MrNood why did you start typing then didn't finish!

OpenStudy (mrnood):

 change in y CP = 10 --2 =12 change in x CP = 10 -1 = 9 so if oyu dilate by 2.5 those values increase by 2.5 so changein in y CP' = 12 * 2.5 = 30 so new point is 10 - 30 0= -20 change in x CP' = 9 *2.5 = 24.5 so new point is 10- 22.5 = -12.5